# Collaborating Vision, Depth, and Thermal Signals for Multi-Modal Tracking: Dataset and Algorithm

### v2

@@ 5.2 Comparative Experiments
We evaluate a range of RGB-only, RGB-D and RGB-T trackers, inclding ATOM danelljan2019atom, DiMP bhat2019learning, PrDiMP danelljan2020probabilistic, ToMP mayer2022transforming, OSTrack ye2022joint, SeqTrack chen2023seqtrack, Mixformer cui2024mixformer, DeT yan2021depthtrack, SPT zhu2023rgbd1k, ViPT zhu2023visual, SDSTrack hou2024sdstrack, TBSI hui2023bridging, BAT cao2024bi, UnTrack wu2024single, and DCEvo liu2025dcevo fusion with OSTrack, and compare them with the proposed RDTTrack. Table 2 reports the tracking performance of various methods on the RGBDT500 test set using AUC and DP metrics. Our RDTTrack achieves 0.752 and 0.792 in AUC and DP, respectively. As shown, the proposed RDTTrack demonstrates superior performance across both metrics, consistently outperforming existing uni-modal and dual-modal trackers. Specifically, among single-modal RGB-only trackers, the MixFormer achieves the best DP and AUC of 0.732 and 0.781. By effectively leveraging additional depth and thermal infrared modalities, RDTTrack achieves superior performance compared to MixFormer, with improvements of 2.0% in AUC and 1.1% in DP, respectively.
For dual-modal tracking approach, while the best RGB-D and RGB-T trackers outperform single-modal approaches, they remain inferior to our tri-modal tracking framework. In detail, ViPT_RGBD, SDSTrack_RGBD and UnTrack_RGBD achieve AUC scores of 0.720, 0.718 and 0.733, and DP scores of 0.759, 0.763 and 0.776, respectively. In comparison, RDTTrack surpasses these two RGB-D tracking approaches with an AUC improvement of 3.2%, 3.4% and 1.9% respectively, and a DP improvement of 3.3%, 2.9% and 1.6%, respectively. Among the RGB-T trackers, the UnTrack_RGBT achieves the optimal AUC of 0.732 and DP of 0.790, outperforming other RGB-T methods like ViPT_RGBT, SDS_RGBT and BAT. In addition, the RGB-T image fusion method DCEvo liu2025dcevo with OSTrack achieves improved performance than RGB-only OSTrack. It clearly validates the effectiveness of image-level fusion of DCEvo for robust multi-modal tracking. The proposed tri-modal tracker, RDTTrack, which incorporates an additional depth modality compared to BAT, achieves performance gains of 3.9% in AUC and 1.0% in DP, respectively. In summary, these results validate the effectiveness of RDTTrack’s architecture in fully exploiting the complementary strengths of RGB, depth, and thermal modalities for enhanced tracking performance.
@@
STARK yan2021learning RGB 0.692 0.732
SPT_RGBD zhu2023rgbd1k RGB+Depth 0.706 0.757
SPT_RGBT zhu2023rgbd1k RGB+Thermal 0.730 0.784
OSTrack ye2022joint RGB 0.694 0.736
ViPT_RGBD zhu2023visual RGB+Depth 0.719 0.762
ViPT_RGBT zhu2023visual RGB+Thermal 0.729 0.768
@@ 5.3 Ablation Studies
In Table 4, we present the performance of SPT and ViPT on the RGBDT500 test set, following their retraining on the RGBDT500 training set. With training on dual-modal RGB-D and RGB-T data of RGBDT500, the SPT and ViPT achieve oblivious performance improvements compared to their single-modal baselines STARK yan2021learning and OSTrack yan2021learning. Nevertheless, our proposed RDTTrack, which integrates RGB, depth, and thermal modalities, consistently outperforms the retrained SPT and ViPT models, demonstrating the advantage of tri-modal fusion.
@@ A.2 Qualitative Analysis
In Fig. 6, we visualized the tracking results of some trackers, including OsTrack (RGB) [ye2022joint], ViPT_RGBD (RGB + Depth), ViPT_RGBT (RGB + Thermal) [zhu2023visual], and RDTTrack (RGB + Depth + Thermal) on several complex frames of RGBDT500. As shown, in challenging scenarios such as low-light environments or modality interference, our RDTTrack effectively leverages the most reliable modality in each scene to achieve more robust and accurate tracking.

### v3

@@ 5.2 Comparative Experiments
We evaluate a range of RGB-only, RGB-D and RGB-T trackers, inclding ATOM Danelljan et al. (2019), DiMP Bhat et al. (2019), PrDiMP Danelljan et al. (2020), ToMP Mayer et al. (2022), OSTrack Ye et al. (2022), SeqTrack Chen et al. (2023), Mixformer Cui et al. (2024), DeT Yan et al. (2021b), SPT Zhu et al. (2023b), ViPT Zhu et al. (2023a), SDSTrack Hou et al. (2024), TBSI Hui et al. (2023), BAT Cao et al. (2024), UnTrack Wu et al. (2024), and DCEvo Liu et al. (2025) fusion with OSTrack, and compare them with the proposed RDTTrack. Table 2 reports the tracking performance of various methods on the RGBDT500 test set using AUC and DP metrics. Our RDTTrack achieves 0.752 and 0.792 in AUC and DP, respectively. As shown, the proposed RDTTrack demonstrates superior performance across both metrics, consistently outperforming existing uni-modal and dual-modal trackers. Specifically, among single-modal RGB-only trackers, the MixFormer achieves the best DP and AUC of 0.732 and 0.781. By effectively leveraging additional depth and thermal infrared modalities, RDTTrack achieves superior performance compared to MixFormer, with improvements of 2.0% in AUC and 1.1% in DP, respectively.
For dual-modal tracking approach, while the best RGB-D and RGB-T trackers outperform single-modal approaches, they remain inferior to our tri-modal tracking framework. In detail, ViPT_RGBD, SDSTrack_RGBD and UnTrack_RGBD achieve AUC scores of 0.720, 0.718 and 0.733, and DP scores of 0.759, 0.763 and 0.776, respectively. In comparison, RDTTrack surpasses these two RGB-D tracking approaches with an AUC improvement of 3.2%, 3.4% and 1.9% respectively, and a DP improvement of 3.3%, 2.9% and 1.6%, respectively. Among the RGB-T trackers, the UnTrack_RGBT achieves the optimal AUC of 0.732 and DP of 0.790, outperforming other RGB-T methods like ViPT_RGBT, SDS_RGBT and BAT. In addition, the RGB-T image fusion method DCEvo Liu et al. (2025) with OSTrack achieves improved performance than RGB-only OSTrack. It clearly validates the effectiveness of image-level fusion of DCEvo for robust multi-modal tracking. The proposed tri-modal tracker, RDTTrack, which incorporates an additional depth modality compared to BAT, achieves performance gains of 3.9% in AUC and 1.0% in DP, respectively. In summary, these results validate the effectiveness of RDTTrack’s architecture in fully exploiting the complementary strengths of RGB, depth, and thermal modalities for enhanced tracking performance.
@@
STARK Yan et al. (2021a) RGB 0.692 0.732
SPT_RGBD Zhu et al. (2023b) RGB+Depth 0.706 0.757
SPT_RGBT Zhu et al. (2023b) RGB+Thermal 0.730 0.784
OSTrack Ye et al. (2022) RGB 0.694 0.736
ViPT_RGBD Zhu et al. (2023a) RGB+Depth 0.719 0.762
ViPT_RGBT Zhu et al. (2023a) RGB+Thermal 0.729 0.768
@@ 5.3 Ablation Studies
In Table 4, we present the performance of SPT and ViPT on the RGBDT500 test set, following their retraining on the RGBDT500 training set. With training on dual-modal RGB-D and RGB-T data of RGBDT500, the SPT and ViPT achieve oblivious performance improvements compared to their single-modal baselines STARK Yan et al. (2021a) and OSTrack Yan et al. (2021a). Nevertheless, our proposed RDTTrack, which integrates RGB, depth, and thermal modalities, consistently outperforms the retrained SPT and ViPT models, demonstrating the advantage of tri-modal fusion.
@@ A.2 Qualitative Analysis
In Fig. 6, we visualized the tracking results of some trackers, including OsTrack (RGB) [52], ViPT_RGBD (RGB + Depth), ViPT_RGBT (RGB + Thermal) [56], and RDTTrack (RGB + Depth + Thermal) on several complex frames of RGBDT500. As shown, in challenging scenarios such as low-light environments or modality interference, our RDTTrack effectively leverages the most reliable modality in each scene to achieve more robust and accurate tracking.